# Fair Ordering via Streaming Social Choice Theory

### v3

@@ Abstract.
Prior work (themis; kelkar2020order; cachin2022quick; vafadar2023condorcet) studies the question of “fairly” ordering transactions in a replicated state machine. Each of $n$ replicas receives transactions over a network in a possibly different order, and the system must aggregate these observed orderings into a single order. We argue that this problem is best viewed through the lens of social choice theory, in which (in the preference aggregation problem) rankings on candidates are aggregated into an election result.
Two features make this problem novel and distinct. First, the number of transactions is unbounded, and an ordering must be defined over a countably infinite set. And second, decisions must be made quickly and with only partial information. Additionally, some faulty replicas might alter their reported observations; their influence on the output should be bounded and well understood.
The prior work studies a “$\gamma$-batch-order-fairness” property, which divides an ordering into contiguous batches. If a $\gamma$ fraction of replicas receive a transaction $\tau$ before another transaction $\tau^{\prime}$, then $\tau^{\prime}$ cannot be in an earlier batch than $\tau$. We strengthen this definition to require that batches have minimal size ($\gamma$-batch-order-fairness can be vacuously satisfied by arbitrarily large batches) while accounting for the possibility of faulty replicas.
The lens of social choice theory allows us to give an ordering protocol with both strictly stronger fairness guarantees and better liveness properties than prior work. We study the Ranked Pairs method. Analysis of how missing information moves through the algorithm allows our streaming version to know when it can output a transaction. Deliberate construction of a tiebreaking rule ensures our algorithm outputs a transaction after a bounded time (in a synchronous network). Prior work relies on a fixed choice of $\gamma$ and bound on the number of faulty replicas $f$, but we show that Ranked Pairs satisfies our definition for every $\frac{1}{2}<\gamma\leq 1$ simultaneously and for any $f$, where fairness guarantees linearly degrade as $f$ increases.
@@ 1. Introduction
We study the problem of ordering transactions in the widely-used replicated state machine architecture (e.g., (ongaro2014search; lamport2001paxos; yin:hotstuff; oki1988viewstamped)). In the standard setting, each of a set of $n$ distinct replicas maintains a copy of a state machine. Replicas communicate to agree on a totally-ordered log of transactions $\tau_{1}\prec\tau_{2}\prec\tau_{3}\prec\dots$, and then each replica applies the transactions in this order to its local copy of the state machine. An (unbounded) set of clients create and broadcast new transactions. A classic example is a banking application, in which the state machine maintains a list of user account balances, and a transaction transfers money from one user to another.
This architecture requires that all replicas apply transactions in the same order. In the banking example, a client $A$ with a \$1 balance might create two transactions, one which sends a \$1 payment to client $B$ and one which sends a \$1 payment to client $C$. Only one of payments can execute successfully (assuming $A$’s balance is not allowed to overdraft) so if different replicas apply these transactions in different orders, then they will disagree on the balances of clients $B$ and $C$.
@@
The key observation for this work is that this problem is a novel streaming variation of the classic preference aggregation problem of social choice theory (hagele2001lulls; llull; colomer2013ramon; condorcet1785essay; arrow1950difficulty). Prior work on this problem (kelkar2020order; themis; cachin2022quick) focuses on network communication, but (through varying protocols) each produces agreement on a “vote” from each replica on the ordering of a set of transactions (in social choice parlance, a “ranking” on a set of “candidates”). Agnostic to the choice of network protocol, how should a system aggregate a set of proposed orders (“rankings”) into a single, total order?
@@
As an additional complication, replicas might adjust their reported orderings in response to strategic considerations. One desirable property of an order aggregation rule is that the influence of a (potentially colluding) subset of replicas is bounded and precisely quantified. Mirroring the prior work, we write here that it is “faulty” behavior for a replica to report anything other than the order in which it observes transactions arriving over the network.
@@ 1.1. Our Results
To the best of our knowledge, this work is the first to directly pose the question as a streaming version of the classic preference aggregation problem. We first observe that for an ordering method to be well-defined on a countably infinite set, it suffices for there to exist a “monotonic” and “asymptotically live” algorithm that implements the ordering method on finite, initial segments of the input. This algorithm must not undo a decision when its (finite) input is extended (”monotonicity”) and must eventually output every transaction (”asymptotic liveness”). The core challenge in building such an algorithm is in determining when an algorithm has enough information to make a decision that is consistent with its hypothetical output on any extension of its input.
@@
As an example, given $f$ faulty replicas out of $n$ total replicas and a fixed $\gamma$, the protocol of kelkar2020order satisfies $(\gamma,\frac{f}{n})$-minimal-batch-batch-order-fairness (Lemma 4.3). Unfortunately, that protocol relies on an explicit choice of a parameter $\gamma$ and an explicit bound on $f$, and it is not at all clear whether a higher or lower parameter $\gamma$ gives a stronger guarantee (§11 gives an example). By contrast, the algorithm we give here achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and for any number of faulty replicas. Fairness guarantees smoothly degrade as the number of faulty replicas increases.  [^1] Our algorithms do not require knowledge of a bound on the number of faulty replicas. This number may be separate from a bound on the number of faulty replicas in an underlying communication protocol. Standard Byzantine-fault tolerant communication protocols require $n\geq 3f+1$.
@@
Analysis of this directed graph gives an important structural lemma about what pieces of information Ranked Pairs uses to determine its output. This lemma allows a streaming version of Ranked Pairs to carefully track how uncertainty about as-yet-unseen transactions propagates throughout Ranked Pairs’s algorithm. Ranked Pairs iterates over edges in order of edge weight. We find that the tiebreaking rule in this ordering is a critical part of the algorithm. Using a naive deterministic rule, our streaming algorithm might never output any transactions. However, we then show in §7 that deliberate construction of the tiebreaking rule guarantees asymptotic liveness.
Finally, a synchronous network assumption combined with our tiebreaking guarantees that our algorithm outputs each transaction after a finite amount of time. If at most $\Delta$ time elapses between when a transaction $\tau$ is sent and when every honest replica votes on it, for example, our algorithm outputs $\tau$ after a delay of at most $O(n\Delta)$. This can be reduced to $O(k\Delta)$, by rounding edge weights to the nearest $\frac{1}{k}$— although this rounding reduces the fairness guarantee to $(\gamma,\frac{f}{n}+\frac{1}{2k})$-minimal-batch-order-fairness (again for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and any $f$).
@@ 2. Preliminaries and System Model
We say that a replica is honest if its true, observed ordering always extends its ordering vote, and if whenever it submits a vote, it contains all transactions that the replica has observed at that time. Otherwise, the replica is faulty. We assume that at most $f$ of the $n$ replicas are faulty, although our algorithms do not require knowledge of $f$. 333 Of course, $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness (Definition 4.4) is only meaningful if $f<\frac{n}{2}$.
In the rest of this work, we assume that every replica eventually votes on every transaction. This may not necessarily hold for faulty replicas. A precise choice of response to this type of faulty behavior will depend on the network model and communication protocol, but, informally, if a replica fails to vote on a transaction within a “sufficient” time, it suffices for the honest replicas to fabricate a vote on behalf of the faulty replica. We address this difficulty in §8.
3. Related Work
3.1. Order-Fairness
Closest to this work are Aequitas (kelkar2020order) and Themis (themis), which propose the definition of $\gamma$-batch-order-fairness and instantiate ordering protocols based on aggregating ordering votes into a total order. Aequitas achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for a fixed choice of $\gamma$, but is not asymptotically live. §4.3 gives a more detailed comparison to these works.
Cachin et al. (cachin2022quick) gives a notion of “$\kappa$-differential-order-fairness;” Theorem D.1 of (themis) shows an equivalence between this definition and $\gamma$-batch-order-fairness, for a particular choice of $\gamma$ (as a function of $\kappa$).
Our definition of $(\gamma,\delta)$-minimal-batch-order-fairness (Definition 4.4) strengthens the (vacuously satisfiable) $\gamma$-batch-order-fairness definition used in prior work. §4 discusses the details of this relationship.
Additionally, Themis’s (themis) round-based sequencing protocol presents a different problem insurmountable by any process like the one suggested by vafadar2023condorcet that sorts batches separately. Namely, Themis may finalize the composition of a batch before has enough information to do guarantee that, according to the Ranked Pairs ordering, all transactions in the batch must come before all those not in the batch (and not already in the output). This means that Themis may output two batches in sequence, such that running Ranked Pairs on each batch separately produces a different output than running Ranked Pairs on the union of the two batches together. We give a concrete example of this phenomenon in §13.
aequitaspermissionless instantiate Aequitas in a setting where the number of replicas is not known.
zhang2020byzantine adds timestamps to ordering votes and sorts transactions by median reported timestamp to provide a linearizability guarantee between transactions. This guarantee is incomparable to $(\gamma,\delta)$-minimal-batch-order-fairness.
3.2. Social Choice
crisman2022voting study the preference aggregation problem where the output must be a cyclic permutation of the candidates.
3.3. Front-Running
The study of order fairness was initially motivated (as in, e.g., kelkar2020order) by the goal of limiting an adversary’s ability to order transactions (“front-running”) in public blockchains for their own financial benefit (daian2020flash). Additional approaches to this front-running problem include commiting to a transaction ordering before any replica can know transaction contents, using threshold encryption or commit-reveal schemes (malkhi2022maximal; clineclockwork; zhang2022flash). Li et. al (li2023transaction) study the problem of verifiably computing an ordering within an encrypted hardware enclave (costan2016intel). These could be applied on top of a streaming ordering algorithm.
Kavousi et. al (kavousi2023blindperm) randomly shuffle blocks of encrypted transactions. Some decentralized exchanges (ramseyer2023speedex; cowswapproblem; penumbraswap) process transactions in unordered batches, thereby eliminating the need to compute an ordering. ferreira2022credible construct an exchange-specific ordering that provides traders with certain worst-case price guarantees. mamageishvili2023buying use a median-timestamp mechanism akin to zhang2020byzantine, but allow clients to pay to reduce its reported timestamp.
@@ 4.1. Batch Order Fairness
Our discussion here follows that of (kelkar2020order; themis; cachin2022quick), but gives, unlike prior work, a definition that is meaningful and achievable.
Honest replicas might receive transactions over the network in a different order. To resolve disagreements “fairly,” recent work (kelkar2020order) proposes a notion known as “$\gamma$-Batch-Order-Fairness,” which we reproduce below from Definition III.1 of themis (a more complicated but functionally similar definition is used in kelkar2020order).
@@
Suppose that $\tau$ and $\tau^{\prime}$ are received by all nodes. If $\gamma n$ nodes received $\tau$ before $\tau^{\prime}$, then a ranking algorithm outputs $\tau$ no later than $\tau^{\prime}$.
Aequitas (kelkar2020order) outputs transactions in batches, and “no later than” means “in the same batch.”Transactions within a batch are ordered arbitrarily. Themis (themis) outputs a total ordering, with the assertion that this total ordering could be divided into contiguous (disjoint) batches satisfying this property.
One challenge is the parameter $\gamma$. It is not obvious if a higher or lower $\gamma$ gives stronger fairness properties. Indeed, there are simple examples (with no faulty replicas) where Definition 4.1 (actually the stronger Definition 4.4) only implies ordering restrictions at high values of $\gamma$, others at only low values, and still others at only intermediate values (§11). Prior work relies on a fixed choice of $\gamma$, but our algorithms satisfy Definition 4.4 for all values of $\gamma$ simultaneously.
@@ 4.2. Batch Minimality
Unfortunately, Definition 4.1 is insufficient to provide meaningful ordering guarantees. A protocol that outputs all transactions in the same batch vacuously satisfies this definition, and if transactions within a batch can be ordered arbitrarily, then every ordering satisfies Definition 4.1 for every $\gamma$.
Prior work (themis) recognizes the need for “minimality”.However, the notion of minimality in prior work is not achievable in the presence of faulty replicas (Theorem 4.2). Towards a feasible notion of minimality, we first define a graph denoting ordering dependencies.
@@
Suppose that replicas report ordering votes $(\sigma_{1},...,\sigma_{n})$ on a set of transactions $V$. The Ordering Graph $G(\sigma_{1},...\sigma_{n})$ is a complete, weighted directed graph with vertex set $V$ and, for each transaction pair $(\tau,\tau^{\prime})$, an edge of weight $w(\tau,\tau^{\prime})=\alpha$ if $\alpha n$ replicas report receiving $\tau$ before $\tau^{\prime}$.
@@
Unfortunately, this notion of minimality is impossible to achieve (for any $\gamma$) in the face of any faulty replicas. The condition on $f$ in Theorem 4.2 is a technical artifact of a construction in the proof (related to the pigeonhole principle). For some parameter settings, such as $n=7$ with $\gamma=\frac{2}{3}$, the condition admits $f=1$.
{theoremrep}
@@
There is one subtle difference between this definition and (an analogue of) the notion of “minimality” in Theorem IV.1 of Themis, and the sequence of batches output in Aequitas. When there are two disjoint strongly connected components with no dependencies (of strength at least $\gamma$) from one to the other, Definition 4.4 allows the output to interleave these components. This may be strictly required (§12).
@@ 4.3. Comparison to Prior Work
Before constructing our protocol, we first apply Definition 4.4 to related prior work.
@@
Aequitas with parameters $(\gamma,f)$ achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness.
@@
By contrast, Ranked Pairs (Theorem 5) achieves $(\gamma,\frac{f}{n})$-minimal-batch-fairness for every $\gamma$ simultaneously, and for any $f$ (our work does not require knowledge of a bound on $f$). Aequitas also is not asymptotically live (Definition 6.2).
Themis claims (Theorem IV.1, (themis)) to output a sequence that can be partitioned into minimal batches, which appears to contradict Theorem 4.2. However, Themis relaxes its notion of an “ordering dependency”.
Where $\gamma$-batch-order-fairness would create a dependency for $\tau$ not after $\tau^{\prime}$ if at least $\gamma n$ replicas receive $\tau$ before $\tau^{\prime}$, the relaxation implied by Definition III.1 of (themis) creates such a dependency if at least $n(1-\gamma)+1$ receive $\tau$ before $\tau^{\prime}$ (that is, it cannot be the case that more than $\gamma n$ receive $\tau^{\prime}$ before $\tau$). Under this definition, the construction in Theorem 4.2 constitutes what the authors of (themis) call a “weak Concorcet Cycle”.
Extra ordering dependencies, however, can artificially merge batches together, giving counterintuitive results. {lemmarep} In Themis (themis), even if all nodes report $\tau$ before $\tau^{\prime}$ and there are no faulty replicas nor cycles in ordering dependencies (as defined in Definition 4.1) for any $\gamma>1/2$, the output may put $\tau^{\prime}$ before $\tau$.
Lemma 4.3 follows from the fact that when two transactions $\tau,\tau^{\prime}$ receive $n-f$ votes in one round of Themis, it determines whether to consider as a dependency $\tau\prec\tau^{\prime}$ or $\tau^{\prime}\prec\tau$ by majority vote (disregarding $\gamma$). The minimality statement of Theorem IV.1 of (themis) is with regard to this weaker notion of dependency.

### v4

# Fair Ordering in Replicated Systems via Streaming Social Choice

@@ Abstract.
How can we order transactions “fairly” in a replicated state machine (of which today’s blockchains are a prototypical example)? In the model of prior work (themis; kelkar2020order; cachin2022quick; vafadar2023condorcet; kiayias2024ordering), each of $n$ replicas observes transactions in a different order, and the system aggregates these observed orderings into a single order. We argue that this problem is best viewed directly through the lens of the classic preference aggregation problem of social choice theory (instead of as a distributed computing problem), in which rankings on candidates are aggregated into an election result.
Two features make this problem novel and distinct. First, the number of transactions is unbounded, and an ordering must be defined over a countably infinite set. And second, decisions must be made quickly and with only partial information. Additionally, some faulty replicas might misreport their observations; the influence of faulty replicas on the output should be well understood.
Prior work studies a “$\gamma$-batch-order-fairness” property, which divides an ordering into contiguous batches. If a $\gamma$ fraction of replicas receive a transaction $\tau$ before another transaction $\tau^{\prime}$, then $\tau^{\prime}$ cannot be in an earlier batch than $\tau$. This definition holds vacuously, so we strengthen it to require that batches have minimal size, while accounting for faulty replicas.
This lens gives a protocol with both strictly stronger fairness and better liveness properties than prior work. We specifically adapt the Ranked Pairs (tideman1987independence) method to this streaming setting. This algorithm can be applied on top of any of the communication protocols (in various network models) of prior work for immediate liveness and fairness improvements. Prior work relies on a fixed choice of $\gamma$ and a bound on the number of faulty replicas $f$, but we show that Ranked Pairs satisfies our definition for every $\frac{1}{2}<\gamma\leq 1$ simultaneously and for any $f$, where fairness guarantees degrade as $f$ increases.
@@ 1. Introduction
We study the problem of ordering transactions in the widely-used replicated state machine architecture (e.g., (ongaro2014search; lamport2001paxos; yin:hotstuff; oki1988viewstamped)). In the standard setting, each of a set of $n$ distinct replicas maintains a copy of a state machine. Replicas communicate to agree on a totally-ordered log of transactions $\tau_{1}\prec\tau_{2}\prec\tau_{3}\prec\dots$, and then each replica applies the transactions in this order to its local copy of the state machine. An (unbounded) set of clients create and broadcast new transactions. This architecture famously underlies many of today’s blockchains, but also underlies many traditional (centralized) services, such as bank infrastructure.
This architecture requires that all replicas apply transactions in the same order. In a bank example, a client $A$ with a \$1 balance might create two transactions, one which sends a \$1 payment to client $B$ and one which sends a \$1 payment to client $C$. Only one of payments can execute successfully (assuming $A$’s balance is not allowed to overdraft) so if different replicas apply these transactions in different orders, then they will disagree on the balances of clients $B$ and $C$.
@@
The key observation for this work is that this problem is a novel streaming variation of the classic preference aggregation problem of social choice theory (hagele2001lulls; llull; colomer2013ramon; condorcet1785essay; arrow1950difficulty). Prior work on this problem (kelkar2020order; themis; cachin2022quick; aequitaspermissionless) use very different network communication protocols, but ultimately each produces agreement on a “vote” from each replica on the ordering of a set of transactions (in social choice parlance, a “ranking” on a set of “candidates”). Agnostic to the choice of network protocol, how should a system aggregate a set of proposed orders (“rankings”) into a single, total order?
@@
One additional consideration is that replicas might strategically adjust their reported orderings. Prior work (kelkar2020order; themis; cachin2022quick) writes that “honest” replicas must report transactions in the order in which they arrive over a network, and any other behavior is “faulty.” One desirable property of an order aggregation rule is that the influence of a (colluding) subset of faulty replicas is precisely bounded.
@@ 1.1. Our Results
We first observe that for an ordering method to be well-defined on a countably infinite set, it suffices for there to exist a “monotonic” and “asymptotically live” algorithm that implements the ordering method on finite, initial segments of the input. This algorithm must not undo a decision when its (finite) input is extended (”monotonicity”) and must eventually output every transaction (”asymptotic liveness”). The core challenge in building such an algorithm is in determining when an algorithm has enough information to make a decision that is consistent with its hypothetical output on any extension of its input.
@@
As an example, given $f$ faulty replicas out of $n$ total replicas and a fixed $\gamma$, the protocol of kelkar2020order satisfies $(\gamma,\frac{f}{n})$-minimal-batch-batch-order-fairness (Lemma 3.3). Unfortunately, that protocol relies on an explicit choice of a parameter $\gamma$ and an explicit bound on $f$, and it is not at all clear whether a higher or lower parameter $\gamma$ gives a stronger guarantee (§10 gives an example). By contrast, the algorithm we give here achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and for any number of faulty replicas. Fairness guarantees smoothly degrade as the number of faulty replicas increases.  [^1] Our algorithms do not require knowledge of a bound on the number of faulty replicas. This notion of “faulty” replica is distinct from notions of “faulty” in classical Byzantine-fault tolerant communication protocols, such as (castro1999practical).
@@
Analysis of this directed graph gives an important structural lemma about what pieces of information Ranked Pairs uses to determine its output. This lemma allows a streaming version of Ranked Pairs to carefully track how uncertainty about as-yet-unseen transactions propagates throughout Ranked Pairs’s algorithm. Ranked Pairs iterates over edges in order of edge weight. An interesting observation is that the tiebreaking rule in this edge ordering critically determines overall liveness. A worst-case (fixed) rule could cause our streaming algorithm to never output any transactions. However, we show in §6 how to construct a tiebreaking rule in a streaming fashion that guarantees not only asymptotic liveness but also an explicitly bounded liveness delay. If, for example, at most $\Delta$ time elapses between when a transaction $\tau$ is sent and when every honest replica votes on it, for example, our algorithm outputs $\tau$ after a delay of at most $O(n\Delta)$. This can be reduced to $O(k\Delta)$, by rounding edge weights to the nearest $\frac{1}{k}$— although this rounding reduces the fairness guarantee to $(\gamma,\frac{f}{n}+\frac{1}{2k})$-minimal-batch-order-fairness (again for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and any $f$).
@@ 2. Preliminaries and System Model
We say that a replica is honest if its true, observed ordering always extends its ordering vote, and if whenever it submits a vote, it contains all transactions that the replica has observed at that time. Otherwise, the replica is faulty. We denote the number of faulty replicas as $f$ (out of $n$), and do not assume a bound or knowledge of $f$. 333 Of course, $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness (Definition 3.4) is only meaningful if $f<\frac{n}{2}$.
In the rest of this work, we assume that every replica eventually votes on every transaction. This may not necessarily hold for faulty replicas. A precise choice of response to this type of faulty behavior will depend on the network model and communication protocol, but, informally, if a replica fails to vote on a transaction within a “sufficient” time, it suffices for the honest replicas to fabricate a vote on behalf of the faulty replica. §16 gives an example construction that satisfies this property, but note that prior work on this problem (e.g., (kelkar2020order; themis; cachin2022quick)) implicitly use the same ideas to handle unresponsive replicas.
@@ 3.1. Batch Order Fairness
Our discussion here follows the notion of “fairness” discussed in prior work (e.g., (kelkar2020order; themis; cachin2022quick; kiayias2024ordering)). The natural notion of “fairness” in prior work, which we follow here, is that if “most” replicas report a transaction $\tau$ before another transaction $\tau^{\prime}$, then the protocol outputs $\tau$ before $\tau^{\prime}$. Making this notion precise requires accounting for what are known as Condorcet cycles. A (super)majority might report $\tau$ before $\tau^{\prime}$, another majority reports $\tau^{\prime}$ before $\tau^{\prime\prime}$, and a third majority reports $\tau^{\prime\prime}$ before $\tau$.
Aequitas (kelkar2020order) sidesteps this problem by outputting these Condorcet cycles in “batches,” which leads to the following definition (paraphrased from (kelkar2020order)).
@@
Suppose that $\tau$ and $\tau^{\prime}$ are received by all nodes. If $\gamma n$ nodes received $\tau$ before $\tau^{\prime}$, then a ranking algorithm never outputs $\tau^{\prime}$ in a later batch than $\tau$.
In other words, $\tau$ and $\tau^{\prime}$ could be in the same batch. Transactions within a batch are ordered arbitrarily. Subsequent work (themis; cachin2022quick; kiayias2024ordering) follows this same definition (or a restatement of it).
There are two key problems with this definition. First, this definition does not preclude a ranking algorithm from putting all transactions into one batch (and therefore ordering transactions arbitrarily). For this definition to be meaningful, it needs a notion of “minimality” to batch sizes. We address this problem by strengthening the definition to include a notion of “minimality” (Definition 3.4). Some of the prior work (kelkar2020order; cachin2022quick; kiayias2024ordering), but not all (themis), implicitly satisfies this definition.
Second, it is not clear whether a higher or lower $\gamma$ gives stronger fairness properties, yet all of the prior work requires a fixed parameter choice. Indeed, there are simple examples (with no faulty replicas) where Definition 3.4 only implies ordering restrictions at high values of $\gamma$, others at only low values, and still others at only intermediate values (§10). It is not the case that satisfying Definition 3.4 for a low value of $\gamma$ implies satisfying the definition for a high value of $\gamma$. We address this problem by constructing a protocol that satisfies Definition 3.4 for all $\frac{1}{2}\leq\gamma\leq 1$ simultaneously.
@@ 3.2. Batch Minimality
To construct a precise definition, we start with a weighted dependency graph over a set of transactions.
@@
Suppose that replicas report ordering votes $(\sigma_{1},...,\sigma_{n})$ on a set of transactions $V$. The Ordering Graph $G(\sigma_{1},...\sigma_{n})$ is a complete, weighted directed graph with vertex set $V$ and, for each transaction pair $(\tau,\tau^{\prime})$, an edge of weight $w(\tau,\tau^{\prime})=\alpha$ if $\alpha n$ replicas report $\tau\prec\tau^{\prime}$.
@@
However, this notion of minimality is impossible to achieve (for any $\gamma$) in the face of any faulty replicas. The condition on $f$ in Lemma 3.2 is a technical artifact of a construction in the proof (related to the pigeonhole principle). For some parameter settings, such as $n=7$ with $\gamma=\frac{2}{3}$, the condition admits $f=1$.
{lemmarep}
@@
There is one subtle difference between this definition and the requirement for an explicit sequence of output batches in prior work (e.g., in (kelkar2020order; themis)). When there are two disjoint strongly connected components with no dependencies (of strength at least $\gamma$) from one to the other, Definition 3.4 allows the output to interleave these components. This may be strictly required (§11).
@@ 3.3. Comparison to Prior Work
Before constructing our protocol, we first apply Definition 3.4 to related prior work. Aequitas (kelkar2020order) is instantiated with a choice of parameter $\gamma$ and a bound $f$ on the number of faulty replicas.
@@
Aequitas (kelkar2020order) with parameters $(\gamma,f)$ achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness.
@@
Broadly speaking, Aequitas, along with the protocols of cachin2022quick and kiayias2024ordering, follow the same pattern. Through (very different) communication protocols, they come to agreement on, for each transaction pair $\tau$ and $\tau^{\prime}$, the fraction of replicas that report $\tau$ before $\tau^{\prime}$. These protocols choose (in advance) a fixed parameter $\gamma$, and then drop from consideration ordering dependences of strength less than $\gamma$. This is equivalent to constructing the ordering dependency graph of Definition 3.2, but then dropping all edges of weight less than $\gamma$ and then forgetting all of the weights on the remaining edges. The strongly connected components of this graph form the batches output in Aequitas.
Aequitas (as well as the protocols of cachin2022quick and kiayias2024ordering) also is not asymptotically live (Definition 5.2). The strongly connected components of this unweighted dependency graph can be arbitrarily large, and these protocols wait until they observe the entirety of a component before outputting any transaction in it. An exception is one version of the protocol in kiayias2024ordering, which adds timestamps to transactions and (given precise network assumptions) can output a component in a streaming fashion.
By contrast, Ranked Pairs (Theorem 4) achieves $(\gamma,\frac{f}{n})$-minimal-batch-fairness for every $\gamma$ simultaneously, and for any $f$ (our work does not require knowledge of a bound on $f$). Our protocol gets these stronger results by using all of the information available in the problem input. Additionally, Ranked Pairs guarantees a bounded liveness delay (although precise end-to-end results depend also on network conditions).
Themis (themis) claims (Theorem IV.1) that its output can be partitioned into “minimal” batches, which appears to contradict Lemma 3.2. However, Themis uses a much weaker notion of “minimal.”
Where $\gamma$-batch-order-fairness would create a dependency for $\tau$ not after $\tau^{\prime}$ if at least $\gamma n$ replicas receive $\tau$ before $\tau^{\prime}$, the relaxation implied by Definition III.1 of (themis) creates such a dependency if at least $n(1-\gamma)+1$ receive $\tau$ before $\tau^{\prime}$ (that is, it cannot be the case that more than $\gamma n$ receive $\tau^{\prime}$ before $\tau$). Theorem IV.1 of (themis) considers its output batches “minimal” if they are singletons when there are no cycles of these dependencies. This notion of minimality is much weaker than what we discuss here, and does not provide any guarantee when there is significant disagreement between replicas on transaction orderings. Lemma 3.3 gives an example highlighting this distinction.
{lemmarep}
In Themis (themis), even if all nodes report $\tau$ before $\tau^{\prime}$ and there are no faulty replicas nor cycles in ordering dependencies (as defined in Definition 3.1) for any $\gamma>1/2$, the output may put $\tau^{\prime}$ before $\tau$.
This kind of counterintuitive behavior is a possible output of Themis’s protocol. When two transactions $\tau,\tau^{\prime}$ receive $n-f$ votes in one round of Themis, it determines whether to consider as an ordering dependency $\tau\prec\tau^{\prime}$ or $\tau^{\prime}\prec\tau$ by majority vote. Equivalently, Themis as a protocol operates in the regime of $\gamma=\frac{1}{2}$, regardless of a chosen parameter value (the choice of $\gamma$ does affect how Themis considers transactions reported by some but not all replicas, which does not affect this discussion).